OpenStudy (anonymous):

how do you find the square roots of -5x^2+ 45=0

OpenStudy (anonymous):

-5x^2+45=0 try to isolate the x^2

OpenStudy (anonymous):

So start with moving the +45 to the 0 side

OpenStudy (anonymous):

so it will look like this

OpenStudy (eamier):

\[\sqrt{45-5x^2}\] like this?

OpenStudy (anonymous):

no u over complicated things

OpenStudy (anonymous):

wait one min

OpenStudy (eamier):

this is the question i think

OpenStudy (anonymous):

wait im scanning the work xD

OpenStudy (eamier):

i forgot equal to zero. then square both side and become like your answer. sorry

OpenStudy (anonymous):

kinda like this

OpenStudy (eamier):

yes

OpenStudy (anonymous):

Then from there -----> -5x^2=-45 Divide by -5

OpenStudy (anonymous):

i got it! ive done that part. how do i finish the problem and find the answer

OpenStudy (anonymous):

okay i'm following

OpenStudy (anonymous):

whemn u divide by -5 u should gey

OpenStudy (anonymous):

get*

OpenStudy (anonymous):

x^2=9

OpenStudy (anonymous):

From there x^2=9 square root it and u should get ----> \[\sqrt{x^2}=\sqrt{9}\]

OpenStudy (anonymous):

thank you i got it!!

OpenStudy (anonymous):

\[\sqrt{x^2}=\sqrt{9} ----> x=3\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

XD Yay! :D

OpenStudy (anonymous):

No problem :D